# Mind the memory: Consistent time reversal removes artefactual scaling of energy dissipation rate and provides more accurate and reliable thermodynamic inference

### v3

@@ II.3 Positivity and behavior for increasing order
III Splitting probabilities in lumped cycle graph $C_{n}$
@@
[…]
@@
III.1 Note on the waiting time distribution
@@ IV Tree graph
IV.1 Construction of tree graph
Consider a tree graph of depth $d$, i.e., vertices are on levels ${U_{0},\ldots,U_{d}}$, where level $i$ contains ${|U_{i}|=2^{i}}$ vertices. For $i<d$, each vertex ${v\in U_{i}}$ has two offsprings ${x,y\in U_{i+1}}$ with rates ${L_{xv}=L_{yv}=\omega/2^{i+1}}$ in $+$ direction and ${L_{vx}=L_{vy}=\kappa/2^{i+1}}$ in -- direction. Vertices in the bottom level $v\in U_{d}$ have edges of rate $\omega/2^{d}$ towards the root, and rate $\kappa/2^{d}$ from the root to each $v$. We refer to Fig. 1a in the Letter. The steady state is the uniform distribution as can be readily seen from the generator $L$.
IV.2 Microscopic dynamics
@@ IV.3 Mistaken dynamics
IV.4 Mesosystem
We now derive the entropy production rate of the lumped mesosystem when memory is accounted for. Via Laplace transforms, we obtain the splitting probabilities. We defer the derivation to the more general setting in Section VII. For the special setting of the tree graph, we obtain using Eq. S69 for the ratio of splitting probabilities
@@ IV.4 Mesosystem
V Sierpinski-type graph
@@
V.1 Decomposition in cycle basis
@@
Figure S3: Basis cycles with non-zero current. These basis elements coincide with the driven cycles of depth $1$ (blue), depth $2$ (green), and depth $3$ (orange) as well as two corrective cycles in each of the outer subgraphs A, B, and C. These correction cycles are indicated by dashed lines in subgraphs $A$.
A driven regular cycle at depth $l$ has an affinity of $A_{l}=3l\log{\omega}$ ). When driven with $\omega=2$, the contribution of each scale is shown in Table. 1.
@@
V.2 Comparison of stochastic estimator to analytic solution
@@
VI Brusselator
For the Brusselator we choose the parameters as in [5]. Note that the state space can be reduced from $500\times 500$ states to e.g. $450\times 450$ states due to the location of the limit cycle in the steady state. This can be verified numerically, by plotting the average time spent in a given state in the steady state. We show this statistic in Fig. S4 for a trajectory of length $5*10^{8}$ after skipping an initial phase of $10^{9}$ steps ensuring convergence to the steady state, where in particular the right and top region is 00 (black). Indeed, summing the steady-state probabilities for $x>450,y>450$, we get $62.45e-$ 6.
@@ VI Brusselator
VII Tree diamond graph
@@ VII.1 Construction
When walking on the tree, we go in positive direction, if we go from a level $U_{i}$ to $U_{i+1\mod 2d+1}$. Hence, the rates of the edges between levels $U_{i}$ and $U_{i+1\mod 2d}$ sum to $\omega$ in $+$ direction and $\kappa$ in -- direction. From $0=Lp$ we see that the steady state is the uniform distribution over all vertices.
@@ VII.2.1 $\dot{S}$ via edges
VII.2.2 Cycle decomposition
Let our diamond tree be the graph $G=(V,E)$. $G$ has a cycle basis $\mathcal{C}=\{c_{1},\ldots,c_{2^{d}}\}$ of size $|\mathcal{C}|=2^{d}$, with each basis cycle $c_{i}$ being the unique cycle of length $12d+$ 1 going through a unique node in level $L_{d}$. To see this, take as spanning tree $T$ the graph $G$ with all edges between level $L_{d}$ and $L_{d+1}$ removed. Adding any single edge between $L_{d}$ and $L_{d+1}$ will create a cycle. We call the added edge a chord and identify uniquely each cycle $c_{i}$ with its chord $h_{i}\in E$. Four of these cycles are highlighted at the central levels in Fig. 5a (inset) of the Letter.
@@ VII.3 Mesosystem
VII.3.1 Splitting probability $\Phi$ for tree lump
@@
[…]
@@
[…]
@@
yielding residue values of $-\frac{\omega}{2\kappa},\frac{\omega}{4\kappa},\frac{\omega}{4\kappa}$ (the first derivative of the denominator is non-zero for all of $s_{1},s_{2},s_{3}$). Thus
@@
since the roots $s_{i}$ have the same values as in the positive direction case. Even stronger, inverting the matrix from Eq. (S61) yields the same denominator for each entry, hence we get the same roots, and hence the same powers of the exponentials in the waiting time distribution. Note that $s_{i}<0$ for all $\lambda,\kappa,\omega>0,i\in[3]$, so that the distributions indeed converge for $t\rightarrow\infty$.
@@
VII.3.2 Splitting probability $\Phi$ for tree diamond lump
@@
VII.4 Average lump size
To get the average lump size, i.e. the average number of vertices per lump, for a given tree diamond graph of depth $d$ with lump depth $l$, we first find the number of tree diamond lumps and tree lumps. For the former, we note that every diamond lump is rooted at level $d-l$, hence the number of tree diamond lumps is
@@ VII.4 Average lump size
For the latter, we observe that $2(2^{d-l}-1)$ many vertices belong to tree lumps, of which each has a size of $12^{l+1}-$ 1, yielding
@@ VII.5 Mistaken dynamics
Suppose the triangle is rooted in level $U_{i}$, with w.l.o.g $i\in\{0,\ldots,d-2l-1\}$. The last level in the upper half of the macroscopic tree diamond that roots a tree lump is level $d-2l-1$. The rate from the triangle upwards, following Eq. (4) in [5], is
@@
with rates for a tree diamond lump rooted at level $i=d-l$ according to [5]
@@
We have $22\cdot\frac{d-l}{l+1}-$ 2 many edge-levels between triangle levels, $2$ edge-levels between triangles and tree-diamond lumps, and one macroscopic edge, leading to an overall mistaken entropy production rate
@@
VIII 6-state model
@@
IX Sufficient condition for semi-Markov process having at most order $2$
@@
X Notion of self-similarity
@@
XI Notion of dissipative cycles
@@
We can determine the set of dissipative cycles via a spanning tree [6]: Let our graph be $G$. We draw below a spanning tree $T$ for the graph of Fig. 5 in the letter in green. The set of all dissipative cycles is spanned by the set of cycles created when closing a single edge $e\in E(G)\setminus E(T)$. We depict in Fig. S7 one such cycle in orange (middle figure) and pink (right figure). Those are the cycles indicated in the inset of the Fig. 5 in the letter.
@@
XII Technicalities of the sampling procedure
XII.1 Exploiting the symmetry of the graph
@@ XII.2 Choice of parameters
For the Brusselator, (Fig. 3b in the Letter), trajectories were drawn after $N_{\rm skip}=10^{9}$ initial steps and are of length $N_{\rm steps}=10^{9}$. The plot shows the average of $25$ trajectories with almost vanishing error bars. We note that the state space is very large with many states having steady-state probability almost 00 but non-vanishing transition rates. In the microscopic setting, $\lambda=1$, this may lead to undersampling of the steady state. Similarly, order $k\geq 4$ would need even longer trajectories.
@@
XII.3 Subsampling for higher orders

### v4

@@ II.3 Positivity and behavior for increasing order
III Apparent Power Law Scaling
Recent works [5, 6, 7, 8] investigated how the energy dissipation rate in self-similar non-equilibrium reaction networks depends on the coarse-graining scale $\lambda\equiv n/n_{s}$ (i.e. the total number of microstates $n$ relative to the number of mesoscopic states $n_{s}$ in a lumped system) in absence of a timescale separation and observed a power-law dependence of the Markovian estimate $\dot{S}_{\rm M}[\hat{\Gamma}^{s}_{t}]\propto\lambda^{-\alpha}$ [5]. These findings were rationalized in terms of an “inverse cascade” of energy dissipation on different scales. In the specific context of active flows generated by the microtubule-kinesin motor system [9], this amounts to the claim that the most energy is spent to generate and maintain the flow at smaller length scales and only a tiny amount at large length scales [5]. While the underlying calculations in Refs. [5, 6, 7, 8] are technically sound within the inconsistent Markovian approximation, their interpretation [5, 7] is physically inconsistent, as we show in the main Letter by means of counterexamples in Fig. 2 that there are no implications between the power-law dependence of the Markovian estimate $\dot{S}_{\rm M}[\hat{\Gamma}^{s}_{t}]\propto\lambda^{-\alpha}$ and the energy dissipated on different length scales.
IV Splitting probabilities in lumped cycle graph $C_{n}$
@@
In Laplace space ($\tilde{p}(s)\equiv\int_{0}^{\infty}p(t)\exp(-st)dt$), we obtain
@@
IV.1 Note on the waiting time distribution
@@ V Tree graph
V.1 Construction of tree graph
Consider a tree graph of depth $d$, i.e., vertices are on levels $U_{0},\ldots,U_{d}$}, where level $i$ contains $|U_{i}|=2^{i}$} vertices. For $i<d$, each vertex $v\in U_{i}$} has two offsprings $x,y\in U_{i+1}$} with rates $L_{xv}=L_{yv}=\omega/2^{i+1}$} in ++ direction and $L_{vx}=L_{vy}=\kappa/2^{i+1}$} in - direction. Vertices in the bottom level $v\in U_{d}$ have edges of rate $\omega/2^{d}$ towards the root, and rate $\kappa/2^{d}$ from the root to each $v$. We refer to Fig. 1a in the Letter. The steady state is the uniform distribution as can be readily seen from the generator $L$.
V.2 Microscopic dynamics
@@ V.3 Mistaken dynamics
V.4 Mesosystem
We now derive the entropy production rate of the lumped mesosystem when memory is accounted for. Via Laplace transforms, we obtain the splitting probabilities. We defer the derivation to the more general setting in Section VIII. For the special setting of the tree graph, we obtain using Eq. S69 for the ratio of splitting probabilities
@@ V.4 Mesosystem
VI Sierpinski-type graph
@@
VI.1 Decomposition in cycle basis
@@
Figure S3: Basis cycles with non-zero current. These basis elements coincide with the driven cycles of depth $1$ (blue), depth $2$ (green), and depth 33 (orange) as well as two corrective cycles in each of the outer subgraphs A, B, and C. These correction cycles are indicated by dashed lines in subgraphs AA.
[…]
@@
VI.2 Comparison of stochastic estimator to analytic solution
@@
VII Brusselator
For the Brusselator we choose the parameters as in [5]. Note that the state space can be reduced from $500\times 500$ states to e.g. $450\times 450$ states due to the location of the limit cycle in the steady state. This can be verified numerically, by plotting the average time spent in a given state in the steady state. We show this statistic in Fig. S4 for a trajectory of length $5*10^{8}$ after skipping an initial phase of $10^{9}$ steps ensuring convergence to the steady state, where in particular the right and top region is 0 (black). Indeed, summing the steady-state probabilities for $x>450,y>450$, we get 2.45e62.45e-6.
@@ VII Brusselator
VIII Tree diamond graph
@@ VIII.1 Construction
When walking on the tree, we go in positive direction, if we go from a level $U_{i}$ to $U_{i+1\mod 2d+1}$. Hence, the rates of the edges between levels $U_{i}$ and $U_{i+1\mod 2d}$ sum to $\omega$ in ++ direction and $\kappa$ in - direction. From $0=Lp$ we see that the steady state is the uniform distribution over all vertices.
@@ VIII.2.1 $\dot{S}$ via edges
VIII.2.2 Cycle decomposition
Let our diamond tree be the graph $G=(V,E)$. $G$ has a cycle basis $\mathcal{C}=\{c_{1},\ldots,c_{2^{d}}\}$ of size $|\mathcal{C}|=2^{d}$, with each basis cycle $c_{i}$ being the unique cycle of length $2d+1$ going through a unique node in level $L_{d}$. To see this, take as spanning tree $T$ the graph $G$ with all edges between level $L_{d}$ and $L_{d+1}$ removed. Adding any single edge between $L_{d}$ and $L_{d+1}$ will create a cycle. We call the added edge a chord and identify uniquely each cycle $c_{i}$ with its chord $h_{i}\in E$. Four of these cycles are highlighted at the central levels in Fig. 5a (inset) of the Letter.
@@ VIII.3 Mesosystem
VIII.3.1 Splitting probability $\Phi$ for tree lump
@@
Since these two states do not influence the probability distribution of the non-absorbing states, we first solve the system of non-absorbing states and then for the absorbing states. The initial distribution is either $\mathds{1}_{W_{0}}$ (in the $\mid+$ setting) or $\mathds{1}_{W_{l}}$ (in the -$\mid$- setting).
To solve the system of $l+1$ differential equations, we Laplace transform ($\tilde{P}_{W_{i}}(s)\equiv\int_{0}^{\infty}P_{W_{i}}(t)\exp(-st)dt$). Solving for the Laplace-transformed distributions reduces to solving the following tridiagonal linear system of equations
@@
where $a\equiv 1$ iff we start at the root (in the $\mid+$ setting) and $a\equiv 0$ else (in the -$\mid$- setting).
For any $l$, we can solve the above equation to obtain $\tilde{P}_{W_{i}}(s)$ explicitly. It remains to perform an inverse Laplace transform, for which we use Cauchy’s residue method. We show the result here for $l=2$ in the $\mid+$ setting, the ”program” being analogous for every other value of $l\in\mathbb{N}$ and directional conditions. Solving Eq. (S61) yields
@@
yielding residue values of $\frac{\omega}{2\kappa},\frac{\omega}{4\kappa},\frac{\omega}{4\kappa}$ (the first derivative of the denominator is non-zero for all of $s_{1},s_{2},s_{3}$). Thus
@@
since the roots $s_{i}$ have the same values as in the positive direction case  [^3] Even stronger, inverting the matrix from Eq. (S61) yields the same denominator for each entry, hence we get the same roots, and hence the same powers of the exponentials in the waiting time distribution.. Note that $s_{i}<0$ for all $\lambda,\kappa,\omega>0,i\in[3]$, so that the distributions indeed converge for $t\rightarrow\infty$.
@@
VIII.3.2 Splitting probability $\Phi$ for tree diamond lump
@@
VIII.4 Average lump size
To get the average lump size, i.e. the average number of vertices per lump, for a given tree diamond graph of depth $d$ with lump depth $l$, we first find the number of tree diamond lumps and tree lumps. For the former, we note that every diamond lump is rooted at level dld-l, hence the number of tree diamond lumps is
@@ VIII.4 Average lump size
For the latter, we observe that $2(2^{d-l}-1)$ many vertices belong to tree lumps, of which each has a size of $2^{l+1}-1$, yielding
@@ VIII.5 Mistaken dynamics
Suppose the triangle is rooted in level $U_{i}$, with w.l.o.g $i\in\{0,\ldots,d-2l-1\}$  [^4] The last level in the upper half of the macroscopic tree diamond that roots a tree lump is level d2l1d-2l-1.. The rate from the triangle upwards, following Eq. (4) in [5], is
@@
with rates for a tree diamond lump rooted at level i=dli=d-l according to [5]
@@
We have $2\cdot\frac{d-l}{l+1}-2$ many edge-levels between triangle levels, $2$ edge-levels between triangles and tree-diamond lumps, and one macroscopic edge, leading to an overall mistaken entropy production rate
@@
IX Interpreting scalings under consistent time reversal
An apparent power-law dependence of $\dot{S}_{\rm M}$ on the coarse-graining scale $\lambda$ generally has no relation with the microscopic dissipation mechanisms. However, meaningful information about the underlying dynamics may still be inferred from coarse-grained data under consistent time reversal. More specifically, if we can find an order $n$ such that $\dot{S}^{\rm est}_{n+k}=\dot{S}^{\rm est}_{n}$ for any $k\geq 1$. If for such $n,k$ we find that $\dot{S}^{\rm est}_{n+k}\propto\lambda^{-\alpha}$ for some $\alpha>0$, one concludes that the coarse graining hides dissipative cycles (this is not true for a scaling of $\dot{S}_{\rm M}$). If in addition, we have information supporting a self-similar network structure, we may also assume that most energy is dissipated on the smallest and less energy on larger scales.
X 6-state model
@@ IX Interpreting scalings under consistent time reversal
For the computation time, we note that this runs even for very high-order estimators on a normal desktop machine. On a Mac Mini computer, calculating all estimators $\dot{S}_{k}^{est}$ (that is, order $1,\ldots,10$) takes on for the red line 15 seconds and for the blue line 25 minutes. Note that this is not on a cluster and not on a GPU, where the algorithm would run even faster. We developed, in addition, a GPU implementation for larger systems.
XI Sufficient condition for semi-Markov process having at most order 22
@@
XII Notion of self-similarity
@@
XIII Comment on Scaling Law under additional time-scale separation Assumption
In Appendix D of the letter, we perform the Gedanken-Experiment of assuming in addition to Ref. [5] time-scale separation. This rectifies the Markovian assumption for scales of the time-scale separation and smaller. We have shown that a scaling of the (observed) entropy production rate is impossible. Here, we give the weights of the associated numerical experiment in Fig. S7.
123456789101112131415161718192021222324252627282930313233343536100110001001100010011000100110001001100010011000100110001001100010011000100110001001100010011000100110001001100010011000100110001001100010011000100110001001100010011000100110001001100010011000100110001001100010011000100110001001100010011000100110001001100010011000100110001001100010011000
Figure S7: Cycle graph $C_{36}$ with directional weights: We have slow edges (blue), fast edges (dark green), and very fast edges (light green). At the lumping of size $6$ (first lump are vertices $1,\ldots,6$, etc.), we observe a time scale separation.
XIV Notion of dissipative cycles
@@
We can determine the set of dissipative cycles via a spanning tree [10]: Let our graph be $G$. We draw below a spanning tree $T$ for the graph of Fig. 5 in the letter in green. The set of all dissipative cycles is spanned by the set of cycles created when closing a single edge $e\in E(G)\setminus E(T)$. We depict in Fig. S8 one such cycle in orange (middle figure) and pink (right figure). Those are the cycles indicated in the inset of Fig. 5 in the letter.
@@
XV Technicalities of the sampling procedure
XV.1 Exploiting the symmetry of the graph
@@ XV.2 Choice of parameters
For the Brusselator, (Fig. 3b in the Letter), trajectories were drawn after $N_{\rm skip}=10^{9}$ initial steps and are of length $N_{\rm steps}=10^{9}$. The plot shows the average of $25$ trajectories with almost vanishing error bars. We note that the state space is very large with many states having steady-state probability almost 0 but non-vanishing transition rates. In the microscopic setting, $\lambda=1$, this may lead to undersampling of the steady state. Similarly, order $k\geq 4$ would need even longer trajectories.
@@
XV.3 Subsampling for higher orders